### v1

@@ 4.3 Clique-sum
In this section, we prove Theorem 5. We begin by proving the following lemma.
@@
Let $H$ be the clique-sum of two graphs, $H_{1}$ and $H_{2}$, with $C$ being the common clique. Let $P$ be an isometric path of $H$ such that $V(P)\subseteq V(H_{1})$, $P$ has an end-vertex in $C$, and $w$ be any vertex of $H_{2}$. Then, the vertices of $P$ can be covered with three $w$-rooted isometric paths in $H$.
@@
Assume for contradiction that there are no three $w$-rooted isometric paths in $H$ that cover all vertices of $V(P)$. Let $u$ and $v$ be the end-vertices of $P$ and $v\in C$. For some isometric path $T$ and two vertices $\{x,y\}\subseteq V(T)$, $T\left(x,y\right)$ shall denote the subpath of $T$ between $x$ and $y$. Let $u_{1}=u$, and for $i\in\{1,2,3\}$, let $Q_{i}$ be a minimal $w$-rooted isometric path that contains $u_{i}$ and covers the maximum number of vertices of $P$, and define $u_{i+1}$ to be the vertex in $V(P)\setminus\displaystyle\bigcup\limits_{j<i+1}V(Q_{j})$ that is at minimum distance from $u$. For $i\in\{1,2,3\}$, $v_{i}\in V(Q_{i})\cap C$ that is at minimum distance from $w$.
Claim 5.17.
@@
Observe that $V(Q_{1})\cap V(P)$ must be a subpath of $P$ and the minimality of $Q_{1}$ implies that the end-vertex of $Q_{1}$ must be in $V(Q_{1})\cap V(P)$. Let $x\neq w$ be an end-vertex of $Q_{1}$ and suppose $x\neq u$. Let $\mathsf{d}\left(v,x\right)=a$ and as $x\neq u$, $\mathsf{d}\left(v,u\right)\geq a+1$. Then $\mathsf{d}\left(v_{1},x\right)\leq a+1$ and furthermore $\mathsf{d}\left(v_{1},u\right)\leq a$. This implies $\mathsf{d}\left(v,u\right)=a+1,\mathsf{d}\left(v_{1},x\right)=a+1$. Now consider the path $T$ induced by $V(Q_{1}\left(w,v_{1}\right))\cup V(P\left(v,x\right))$. Clearly, $T$ is an $(w,x)$-isometric path ad together with $Q_{1}\left(w,u\right)$ covers all vertices of $P$, which contradicts our assumption that $P$ cannot be covered with three $w$-rooted paths. ∎
Similar arguments, as used in the proof of 5.17, imply the following.
Claim 5.18.
@@
For $i\in\{1,2,3\}$, let $u^{\prime}_{i}$ be the vertex in $V(P)\cap V(Q_{i})$ which is adjacent to $u_{i+1}$. Let $\mathsf{d}\left(u^{\prime}_{1},v\right)=a$, $\mathsf{d}\left(w,v_{1}\right)=b$, and $\mathsf{d}\left(u^{\prime}_{2},u_{2}\right)=\ell$. 5.18 implies that the vertices $u_{1}=u,u^{\prime}_{1},u_{2},u^{\prime}_{2},u_{3},u^{\prime}_{3},u_{4}$ appear in that order when $P$ is traversed from $u$ to $v$.
Claim 5.19.
@@
Since $\mathsf{d}\left(v,u^{\prime}_{1}\right)=a$ and $v$ is adjacent to $v_{1}$, we have $a-1\leq\mathsf{d}\left(u^{\prime}_{1},v_{1}\right)\leq a+1$. If $\mathsf{d}\left(v_{1},u^{\prime}_{1}\right)=a+1$, then $V(Q_{1}\left(w,v_{1}\right))\cup V(P)$ induces a $(w,u_{1})$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, which is a contradiction. Now assume, $\mathsf{d}\left(v_{1},u^{\prime}_{1}\right)=a$. observe that the path given by $V(P\left(u_{2},v\right))\cup V(Q_{1}\left(v_{1},w\right))$ cannot be isometric (otherwise, two $w$-rooted isometric paths cover all vertices of $P$). Hence, $\mathsf{d}\left(u_{2},w\right)\leq a+c-1$ and let $Z$ be any $(w,u_{2})$-isometric path. observe that $Z\cup\left\{u^{\prime}_{1}\right\}$ is a $(w,u^{\prime}_{1})$-isometric path and therefore, $V(Z)\cup V(P\left(u_{2},u\right))$ is an $(w,u)$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, a contradiction. ∎
Claim 5.20.
@@
5.19 implies that $a+b-2\leq\mathsf{d}\left(w,u_{2}\right)\geq a+b$. If $\mathsf{d}\left(w,u_{2}\right)=a+b-2$, then any $(w,u_{2})$-isometric path concatenated with $P\left(u_{2},u\right)$ yields a $(w,u)$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, which is a contradiction. If $\mathsf{d}\left(u_{2},w\right)=a+b$, then $V(P\left(u_{2},v\right))\cup V(Q_{1}\left(v_{1},w\right))$ yields an isometric path that covers strictly more vertices of $P\left(u_{2},v\right)$ than $Q_{1}$, a contradiction. ∎
@@
Observe that $\mathsf{d}\left(v,u_{3}\right)=a-1-\ell$. If $\mathsf{d}\left(u_{3},w\right)=a+b-\ell$, then $V(Q_{1}\left(v_{1},w\right))\cup V(P\left(v,u_{3}\right))$ yields an isometric path that covers strictly more vertices of $P\left(u_{3},v\right)$ than $Q_{3}$, a contradiction. If $\mathsf{d}\left(u_{3},w\right)=a+b-\ell-2$, then 5.20 implies that $V(Q_{3}\left(w,u_{3}\right))\cup V(P\left(u^{\prime}_{2},u_{2}\right))$ induces an $(w,u_{2})$-isometric path that covers strictly more vertices than $Q_{2}$, a contradiction. ∎
@@
Let $G_{1},G_{2},\ldots,G_{t}$ be graphs, each with strong isometric path complexity at most $k$ and $G=\displaystyle\bigoplus\limits_{i=1}^{t}G_{i}$. Let $P$ be any isometric path of $G$ and $v$ be any vertex of $G$. We shall show that the vertices of $P$ can be covered with at most $183k+$ 18 many $v$-rooted isometric paths. First, we make the following observations.
Observation 2.
For a subset $X\subseteq[t]$, $\bigoplus\limits_{i\in X}G_{i}$ is an isometric subgraph of $G$.
Observation 3.
For an integer $i\in[t]$, $V(P)\cap V(G_{i})$ is either empty or induces a subpath of $P$.
@@
Let $w$ be a vertex such that there exists an integer $i\in[t]$ such that $V(P)\cap V(G_{i})$ is non-empty and $w\in V(G_{i})$. Then, the vertices of $P$ can be covered with $k+6$ many $w$-rooted isometric paths in $G$.
@@
3 implies that $V(P)\cap V(G_{i})$ induces a subpath, say $P_{i}$, of $P$. Let $Q$ be a connected component of $V(P)\setminus V(P_{i})$. observe that there exists a clique $C$ in $G$ that separates $Q$ from $V(G_{i})\setminus V(C)$. Let $H_{1}$ and $H_{2}$ be the two connected components in $V(G)\setminus C$ that contain $Q$ and $V(G_{i})\setminus V(C)$, respectively. Let $H$ be the clique-sum of $H_{1},H_{2}$ and $C$ be the common clique. By 2, $H$ is an isometric subgraph of $G$.
Observe that one end-vertex of $Q$ is adjacent to exactly one vertex of $C$, say $u$, of $P_{i}$. Let $Q^{+}$ denote the path induced by $V(Q)\cup\{u\}$. Observe that $Q^{+}$ is an isometric path of $H$, $V(Q^{+})\subseteq V(H_{1})$, and $w$ is a vertex of $H_{2}$. Now, Lemma 16 implies that the vertices of $Q^{+}$ can be covered with three $w$-rooted isometric paths of $H$.
To complete the proof of the lemma, observe that there can be at most two connected components of $V(P)\setminus V(G_{i})$. The above arguments imply that all vertices of $V(P)\setminus V(P_{i})$ can be covered with six $w$-rooted isometric paths in $G$. Since $sipco\left(G_{i}\right)\leq k$, the vertices of $P_{i}$ can be covered with at most $k$ many $w$-rooted isometric paths of $G$. Hence, $P$ can be covered with $k+6$ many $w$-rooted isometric paths of $G$. ∎
Completion of the proof of the theorem: Observe that if there exists an integer $i\in[t]$ such that $V(P)\cap V(G_{i})$ is non-empty and $v\in V(G_{i})$, then the theorem follows from Lemma 17. Otherwise, there exists an integer $i\in[t]$ such that $V(G_{i})\cap V(P)$ is non-empty and there exists a clique $C$ such that $v$ and $P$ lies in different connected components in the graph induced by $V(G)\setminus C$. Let $v^{\prime}$ be any vertex of $C$. Lemma 17 implies that there is a set $\mathcal{P}$ of $k+6$ many $v^{\prime}$-rooted isometric paths in $G$ such that all vertices of $P$ are covered by the paths in $\mathcal{P}$. Now, by Lemma 16, each path in $\mathcal{P}$ can be covered with three $v$-rooted isometric paths in $G$. Therefore, the vertices of $P$ can be covered with $183k+$ 18 many $v$-rooted isometric paths in $G$.
@@ 5 Conclusion
We proved that $U_{t}$-asymptotic minor-free graphs have bounded strong isometric path complexity. However, it is easy to construct a graph class whose strong isometric path complexity is small but contains all $U_{t}$-asymptotic minors. For $t\geq 1$, consider $K_{1,t}$ and replace its edges with a path of length $t$. Let $u_{1},u_{2},\ldots,u_{t}$ denote the leaves. Let $H_{t}$ denote the graph obtained by introducing a path of length $t$ between $u_{i}$ and $u_{i+1}$ for each $i\in[t-1]$. Observe that $sipco\left(H_{t}\right)\leq 2$, but =$\mathcal{H}=\{H_{t}\}_{t\in\mathbb{N}}$ contains $U_{t}$-asymptotic minors for all $t$. Therefore, $U_{t}$-asymptotic minors do not provide a structural characterization of graphs with bounded strong isometric path complexity. It would be interesting to obtain such a characterization.
We proved that important graph operations like fixed powers, line graphs, and clique-sum preserve strong isometric path complexity. However, it is not clear if our upper bounds are tight. It would also be interesting to characterize which graph operations preserve strong isometric path complexity.
Graphs with bounded strong isometric path complexity generalize both hyperbolic graphs and $U_{t}$-asymptotic minor-free graphs and their subclasses. They can be seen as an extension of hyperbolic graphs, where arbitrarily large isometric cycles are allowed to exist (although in a restricted fashion).
It is more generally of interest to see which interesting properties can be lifted from them upto graphs of bounded strong isometric path complexity. This could shed new light on these classes.
It would also be of interest to discover other algorithmic applications of (strong) isometric path complexity beyond Isometric Path Cover.

### v2

@@ 4.3 Clique-sum
In this section, we prove Theorem˜9. We begin by proving the following lemma.
@@
Let $H$ be the clique-sum of two graphs, $H_{1}$ and $H_{2}$, with $C$ being the common clique. Let $P$ be an isometric path of $H$ such that $V(P)\subseteq V(H_{1})$, $P$ has an end-vertex in $C$, and $w$ be any vertex of $H_{2}$. Then, the vertices of $P$ can be covered by three $w$-rooted isometric paths in $H$.
@@
By definition $C\subseteq V(H_{1})$ and $C\subseteq V(H_{2})$. Assume for contradiction that there are no three $w$-rooted isometric paths in $H$ that cover all vertices of $V(P)$. Let $u$ and $v$ be the end-vertices of $P$ and $v\in C$. For some isometric path $T$ and two vertices $\{x,y\}\subseteq V(T)$, $T\left(x,y\right)$ shall denote the subpath of $T$ between $x$ and $y$. Let $u_{1}=u$, and for $i\in\{1,2,3\}$, let $Q_{i}$ be a vertex-minimal $w$-rooted isometric path that contains $u_{i}$ and covers the maximum number of vertices of $P$, and define $u_{i+1}$ to be the vertex in $V(P)\setminus\displaystyle\bigcup\limits_{j\leq i}V(Q_{j})$ that is at minimum distance from $u$. For $i\in\{1,2,3,4\}$, let $v_{i}\in V(Q_{i})\cap C$ that is at minimum distance from $w$. See Figure˜11 for illustration of the notations.
[…]
Figure 11: Notations used in the proof of Lemma˜29.
Claim 4.3.Q.
@@
Observe that $V(Q_{1})\cap V(P)$ must be a subpath of $P$, and the minimality of $Q_{1}$ implies that the end-vertex of $Q_{1}$ distinct from $w$ must be in $V(Q_{1})\cap V(P)$. Let $x$ be the end-vertex of $Q_{1}$ which is distinct from $w$, and suppose that $x\neq u$. Let $\mathsf{d}\left(v,x\right)=a$ and as $x\neq u$, $\mathsf{d}\left(v,u\right)\geq a+1$. Then, $\mathsf{d}\left(v_{1},x\right)\leq a+1$. Since $u$ lies in an isometric path between $v_{1}$ and $x$, it follows that $\mathsf{d}\left(v_{1},u\right)\leq a$. This implies that $\mathsf{d}\left(v,u\right)=a+1$ and $\mathsf{d}\left(v_{1},x\right)=a+1$. Since $\mathsf{d}\left(v,x\right)=a$ (by assumption), the path induced by $\{v_{1}\}\cup V(P\left(v,x\right))$ also has length $a+1$ and is therefore a $(v_{1},x)$-isometric path. Now, consider the path $T$ induced by $V(Q_{1}\left(w,v_{1}\right))\cup V(P\left(v,x\right))$. If $T$ is a $(w,x)$-isometric path, then, together with $Q_{1}\left(w,u\right)$, it covers all vertices of $P$, which contradicts our assumption that $P$ cannot be covered by three $w$-rooted paths. Otherwise, it must be the case that $T\left(w,v\right)$ is not isometric and, therefore, $\mathsf{d}\left(w,v\right)\leq\mathsf{d}\left(w,v_{1}\right)$. But in this case, the path obtained by appending a $(w,v)$-isometric path to $P\left(v,u_{1}\right)$ provides the necessary contradiction. ∎
Similar arguments, as used in the proof of ˜4.3.Q, imply the following.
Claim 4.3.R.
@@
For $i\in\{1,2,3\}$, let $u^{\prime}_{i}$ be the vertex in $V(P)\cap V(Q_{i})$ which is adjacent to $u_{i+1}$. Let $\mathsf{d}\left(u^{\prime}_{1},v\right)=a$, $\mathsf{d}\left(w,v_{1}\right)=b$, and $\mathsf{d}\left(u^{\prime}_{2},u_{2}\right)=\ell$. ˜4.3.R implies that the vertices $u_{1}=u,u^{\prime}_{1},u_{2},u^{\prime}_{2},u_{3},u^{\prime}_{3},u_{4}$ appear in that order when $P$ is traversed from $u$ to $v$.
Claim 4.3.S.
@@
Since $\mathsf{d}\left(v,u^{\prime}_{1}\right)=a$ and $v$ is adjacent to $v_{1}$, we have $a-1\leq\mathsf{d}\left(u^{\prime}_{1},v_{1}\right)\leq a+1$.
Suppose $\mathsf{d}\left(v_{1},u^{\prime}_{1}\right)=a+1$. If $\mathsf{d}\left(w,v\right)\leq\mathsf{d}\left(w,v_{1}\right)$, then the path obtained by appending a $(w,v)$-isometric path with $P\left(v,u_{1}\right)$ is isometric and along with $Q_{1}$ covers $P$, which is a contradiction. Hence, $\mathsf{d}\left(w,v\right)>\mathsf{d}\left(w,v_{1}\right)$. In this case, the vertices $V(Q_{1}\left(w,v_{1}\right))\cup V(P)$ induces a $(w,u_{1})$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, which is a contradiction.
Now assume that, $\mathsf{d}\left(v_{1},u^{\prime}_{1}\right)=a$. Observe that the path given by $V(P\left(u_{2},v\right))\cup V(Q_{1}\left(v_{1},w\right))$ cannot be isometric (otherwise two $w$-rooted isometric paths cover all vertices of $P$). Hence, $\mathsf{d}\left(u_{2},w\right)\leq a+b-1$ and let $Z$ be any $(w,u_{2})$-isometric path. Observe that $Z\cup\left\{u^{\prime}_{1}\right\}$ is a $(w,u^{\prime}_{1})$-isometric path and therefore, $V(Z)\cup V(P\left(u_{2},u\right))$ is an $(w,u)$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, a contradiction. ∎
Claim 4.3.T.
@@
˜4.3.S implies that $a+b-2\leq\mathsf{d}\left(w,u_{2}\right)\leq a+b$. If $\mathsf{d}\left(w,u_{2}\right)=a+b-2$, then any $(w,u_{2})$-isometric path concatenated with $P\left(u_{2},u\right)$ yields a $(w,u)$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, which is a contradiction. If $\mathsf{d}\left(u_{2},w\right)=a+b$, then $V(P\left(u_{2},v\right))\cup V(Q_{1}\left(v_{1},w\right))$ yields an isometric path that covers strictly more vertices of $P\left(u_{2},v\right)$ than $Q_{1}$, a contradiction. ∎
@@
Observe that $\mathsf{d}\left(u_{3},w\right)\leq a+b-\ell-1$ follows from the sentence before the claim. If $\mathsf{d}\left(u_{3},w\right)=a+b$-$\ell$, then $V(Q_{1}\left(v_{1},w\right))\cup V(P\left(v,u_{3}\right))$ yields an isometric path that covers strictly more vertices of $P\left(u_{3},v\right)$ than $Q_{3}$, a contradiction. If $\mathsf{d}\left(u_{3},w\right)=a+b-\ell-2$, then ˜4.3.T implies that $V(Q_{3}\left(w,u_{3}\right))\cup V(P\left(u^{\prime}_{2},u_{2}\right))$ induces an $(w,u_{2})$-isometric path that covers strictly more vertices than $Q_{2}$, a contradiction. ∎
@@
Let $G_{1},G_{2},\ldots,G_{t}$ be graphs, each with strong isometric path complexity at most $k$ and $G=\displaystyle\bigoplus\limits_{i=1}^{t}G_{i}$. Let $P$ be any isometric path of $G$ and $v$ be any vertex of $G$. We shall show that the vertices of $P$ can be covered by at most $3k+18$ many $v$-rooted isometric paths. First, we make the following observations.
Observation 30.
For a subset $X\subseteq[t]$, connected components of $\bigoplus\limits_{i\in X}G_{i}$ are isometric subgraphs of $G$.
Observation 31.
For any integer $i\in[t]$, $V(P)\cap V(G_{i})$ is either empty or induces a subpath of $P$.
@@
Let $w$ be a vertex such that there exists an integer $i\in[t]$ such that $V(P)\cap V(G_{i})$ is non-empty and $w\in V(G_{i})$. Then, the vertices of $P$ can be covered by $k+6$ many $w$-rooted isometric paths in $G$.
@@
˜31 implies that $V(P)\cap V(G_{i})$ induces a subpath, say $P_{i}$, of $P$. Let $Q$ be a connected component of $V(P)\setminus V(P_{i})$. Observe that there exists a clique $C$ in $G$ that separates $Q$ from $V(G_{i})\setminus V(C)$. Let $H_{1}$ and $H_{2}$ be the two connected components in $V(G)\setminus C$ that contain $Q$ and $V(G_{i})\setminus V(C)$, respectively. Let $H$ be the clique-sum of the graphs induced by $V(H_{1})\cup C$ and $V(H_{2})\cup C$. Hence, $C$ is the common clique. By ˜30, $H$ is an isometric subgraph of $G$.
Observe that one end-vertex of $Q$ is adjacent to exactly one vertex of $C$, say $u$, of $P_{i}$. Let $Q^{+}$ denote the path induced by $V(Q)\cup\{u\}$. Observe that $Q^{+}$ is an isometric path of $H$, $V(Q^{+})\subseteq V(H_{1})$, and $w$ is a vertex of $H_{2}$. Now, Lemma˜29 implies that the vertices of $Q^{+}$ can be covered by three $w$-rooted isometric paths of $H$.
To complete the proof of the lemma, observe that there can be at most two connected components of $V(P)\setminus V(G_{i})$. The above arguments imply that all vertices of $V(P)\setminus V(P_{i})$ can be covered by six $w$-rooted isometric paths in $G$. Since $sipco\left(G_{i}\right)\leq k$, the vertices of $P_{i}$ can be covered by at most $k$ many $w$-rooted isometric paths of $G$. Hence, $P$ can be covered by $k+6$ many $w$-rooted isometric paths of $G$. ∎
Completion of the proof of Theorem˜9. Observe that if there exists an integer $i\in[t]$ such that $V(P)\cap V(G_{i})$ is non-empty and $v\in V(G_{i})$, then the theorem follows from Lemma˜32. Otherwise, there exists an integer $i\in[t]$ such that $V(G_{i})\cap V(P)$ is non-empty and there exists a clique $C$ such that $v$ and $P$ lies in different connected components in the graph induced by $V(G)\setminus C$. Let $v^{\prime}$ be any vertex of $C$. Lemma˜32 implies that there is a set $\mathcal{P}$ of $k+6$ many $v^{\prime}$-rooted isometric paths in $G$ such that all vertices of $P$ are covered by the paths in $\mathcal{P}$. Now, by Lemma˜29, each path in $\mathcal{P}$ can be covered by three $v$-rooted isometric paths in $G$. Therefore, the vertices of $P$ can be covered by $3k+18$ many $v$-rooted isometric paths in $G$.
Remark 2.
A possible approach to reduce the constants in Theorem˜9 would be to consider an integer $i$ such that $V(G_{i})\cap V(P)$ is nonempty, a clique $C$ in $G_{i}$ such that $P$ and $v$ lie in different components of $V(G)\setminus C$, and a vertex $w\in C$ which is closest to $v$ in $G$. Then, one might be tempted to consider $k+6$ many $w$-rooted paths that cover $P$ and then append them with $(w,v)$-isometric paths in $G$. However, the resulting paths may not be isometric in $G$, and therefore the above natural approach does not seem to work directly.
@@ 5 Conclusion
We proved that $U_{t}$-asymptotic minor-free graphs have bounded strong isometric path complexity. On the other hand, it follows from [17] that for $k\geq 3$, $(2\times k)$-grid-asymptotic minor-free graphs have unbounded (strong) isometric path complexity. However, the construction of [17] contains vertices of large degree. On the other hand, the graphs constructed to prove Theorem˜7 have bounded maximum degree but do not yield $(2\times k)$-grid-asymptotic minor-free graphs for any constant $k$. We conjecture the following.
Conjecture 33.
There is a function $f\colon\mathbb{N}\times\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$ such that the graphs with maximum degree $d$ and excluding $(2\times t)$-grids as a $K$-fat minor have (strong) isometric path complexity at most $f(d,t,K)$.
We proved that monoholed graphs are $U_{4}$-asymptotic minor-free, which is of independent interest. Thus, these graphs have bounded strong isometric path complexity. For an integer $\ell\geq 1$, let us call a graph $\ell$-polyholed if the total number of distinct lengths of the induced cycles in $G$ is at most $\ell$. We ask the following question.
Question 34.
Is there a function $f\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $\ell$-polyholed graphs with $\ell\geq 1$ have (strong) isometric path complexity at most $f(\ell)$?
It is easy to construct a graph class whose strong isometric path complexity is small but contains all $U_{t}$-asymptotic minors. For $t\geq 1$, consider $K_{1,t}$ and replace its edges with a path of length $t$. Let $u_{1},u_{2},\ldots,u_{t}$ denote the leaves. Let $H_{t}$ denote the graph obtained by introducing a path of length $t$ between $u_{i}$ and $u_{i+1}$ for each $i\in[t-1]$. Observe that $sipco\left(H_{t}\right)\leq 2$, but $\mathcal{H}=\{H_{t}\}_{t\in\mathbb{N}}$ contains $U_{t}$-asymptotic minors for all $t$. Therefore, $U_{t}$-asymptotic minors do not provide a structural characterization of graphs with bounded strong isometric path complexity. Since the property of bounded (strong) isometric path complexity is not preserved under quasi-isometry, we do not expect a coarse characterisation of graphs with bounded strong isometric path complexity. Nevertheless, it would be interesting to obtain a metric characterisation of graphs with bounded strong isometric path complexity. It would also be interesting to characterize graphs with small values of strong isometric path complexity.
We proved that important graph operations like fixed powers, line graphs, and clique-sum preserve strong isometric path complexity. However, it is not clear whether our upper bounds are tight. It would also be interesting to characterize which other graph operations preserve strong isometric path complexity.
Graphs with bounded strong isometric path complexity generalize both hyperbolic graphs and $U_{t}$-asymptotic minor-free graphs and their subclasses. They can be seen as an extension of hyperbolic graphs, where arbitrarily large isometric cycles are allowed to exist (although in a restricted fashion). It is more generally of interest to see which interesting properties can be lifted from them upto graphs of bounded strong isometric path complexity. This could shed new light on these classes. It would also be of interest to discover more algorithmic applications of (strong) isometric path complexity.